OpenStudy (anonymous):

show that e power x is greater than 1+x

OpenStudy (anonymous):

\[e ^{x}=1+\frac{ x }{ 1! }+\frac{ x ^{2} }{ 2! }+\frac{ x ^{3} }{3! }+........>1+x,if x>0\]